# DISL: Fueling Research with A Large Dataset of Solidity Smart Contracts

### v1

@@ 1 Introduction
Smart contracts are self-executing programs running on top of a decentralized ledger [23]. Their use-cases range from financial applications, governance to art. Solidity stands as one of the most used programming languages for smart contract development.
Security and safety of smart contracts is crucial because their deployment on an immutable ledger makes vulnerabilities and bugs irrevocable, risking significant financial losses. Thus, it is essential to conduct thorough testing, analysis, and repair of buggy smart contract code. Tools and techniques employed for this purpose range from static to dynamic analysis [16, 19]. The performance of these tools however is overshadowed by the number of recent successful attacks performed on real projects deployed on chain  [11]. Recent studies indicate that the reliance on synthetic, small, and non-diverse smart contract datasets fail to represent the complexity of real-world projects. Those small datasets hinder the development and assessment of good tools, a core reason for their ineffectiveness [17, 21, 16].
A dataset of real-world smart contracts would aid researchers and developers in different tasks, incl. the production of smart contract development tools. Moreover, as AI-assisted tools gain popularity in smart contract analysis, repair, and synthesis tasks, the demand for large datasets for sake of machine learning tasks on contracts grows increasingly critical [14, 24, 13].

### v2

@@ 1 Introduction
Smart contracts are self-executing programs running on top of a decentralized ledger [23]. Their use-cases range from financial applications, governance, gaming to art. Solidity stands as one of the most used programming languages for smart contract development.
Security and safety of smart contracts is crucial because their deployment on an immutable ledger makes vulnerabilities and bugs irrevocable, risking significant financial losses. Thus, it is essential to conduct thorough testing, analysis, and repair of buggy smart contract code. Tools and techniques employed for this purpose use both static to dynamic analysis [16, 19]. The performance of these tools however is overshadowed by the number of recent successful attacks performed on real projects deployed on chain [11]. Recent studies indicate that the reliance on synthetic, small, and non-diverse smart contract datasets fail to represent the complexity of real-world projects. Those small datasets hinder the development and assessment of reliable tools, a core reason for their ineffectiveness [17, 21, 16].
A dataset of real-world smart contracts will aid researchers and developers in different tasks, including the production of smart contract development tools. Moreover, as AI-assisted tools gain popularity in smart contract analysis, repair, and synthesis tasks, the demand for large datasets for sake of machine learning tasks on contracts grows increasingly critical [14, 24, 13].